OpenStudy (anonymous):

Solve the following system of equations by graphing. x + y = 8 x - 2y = -4 What is the y-value of their intersection? -4 4 12 none of the above

OpenStudy (anonymous):

this is what you are looking for http://www.wolframalpha.com/input/?i=solve++x+%2B+y+%3D+8+and+x+-+2y+%3D+-4+

OpenStudy (anonymous):

Oh heck no, the answer is Kush

OpenStudy (anonymous):



OpenStudy (anonymous):

i got -4?? hmmm

OpenStudy (anonymous):

and I got kush.

OpenStudy (anonymous):

what the heck isn't that w****

OpenStudy (anonymous):

lets check .. \[x + y = 8\] \[x = 8 - y \] replace x = 8 - y in the equation : x - 2y = -8 \[8 - y - 2y = -4\] \[-3y = -12\] \[y = \frac{ 12 }{ 3 }\] \[y = 4\] ( the y-value of intersection )

OpenStudy (anonymous):

that makes sense thank you